# The Limited Impact of Medical Adaptation of Large Language and Vision-Language Models

### v1

@@ 6 Discussion and Conclusion
In this work, we investigated the effectiveness of DAPT for training medically specialized LLMs and autoregressive VLMs suitable for various medical (visual) QA tasks. To that end, we compared several pairs of state-of-the-art medical LLMs/VLMs to their general-domain counterparts, whose only differences lie in medical DAPT and are exactly identical in model architecture and scale. Our work diverges from prior works by providing a direct apples-to-apples comparison of medical and general-domain models while accounting for LLM/VLM sensitivity to prompting details and assessing the statistical significance of the results. We also ensure that our evaluations are comprehensive, by (i) comparing the models in both the zero-/few-shot prompting and supervised fine-tuning regimes, and (ii) covering both QA tasks focused on assessing medical knowledge and those based on real-world clinical notes.
@@
In the zero-/few-shot prompting regime, we found that across both model classes and all model scales, the performance benefits from medical DAPT largely disappear when we (i) tailor the prompt format and choice of few-shot examples to each medical and general-domain model separately; and (ii) account for statistical uncertainty in model comparison (Section 4.1). In particular, we found that when we optimize the prompt only for the medical model and compare each model pair based on their absolute accuracies without accounting for uncertainty, the performance improvements from medical DAPT can be overestimated (Section 4.2), potentially leading to unreliable conclusions about the benefits of medical DAPT on zero-/few-shot prompting performance.
@@
In the supervised fine-tuning regime, we found that medical LLMs overall do show statistically significant improvements on textual QA tasks focused on assessing medical knowledge (Section 5) but not on QA tasks that based on real-world clinical notes (Section 5.2). For medical VLMs, we found that they overall fail to show significant improvements on the visual medical QA tasks that we consider, which we present with the caveat that our findings are based on a relatively small number of VLMs and datasets (Section 5.3).
@@
Our findings suggest that for state-of-the-art general-domain LLMs and VLMs, the performance benefits from additionally pretraining on medical data from public sources such as PubMed may be limited. In fact, almost all of the medical models used in our evaluation use PubMed as the primary source of pretraining data for medical adaptation (Table 1), while open-source datasets commonly used for pretraining the general-domain base models in the first place (e.g., the Pile (Gao et al., 2020), S2ORC (Lo et al., 2020)) often already include PubMed data. Thus, we argue that any claims about improvement from a proposed medical DAPT procedure should be evidenced by rigorous head-to-head comparisons against the corresponding general-domain model to draw reliable conclusions about its effectiveness.
7 Limitations
@@
We discuss our findings with the following caveats. First, there is a vast and growing set of papers on applying medical DAPT to various general-domain base models, and we could not hope to compare all publicly available models here. While we selected the models to cover a wide range of general-domain base models and model scales (7B–70B) (Table 1) and included some of the latest models (e.g., OpenBioLLM and Llama-3), it is always possible that some newly released models do in fact yield better zero-/few-shot or supervised fine-tuning performance on medical QA.
@@
Third, due to computational constraints, we do not consider full fine-tuning of models and exclude the larger 70B-parameters from evaluation, for evaluations in the supervised fine-tuning regime. Prior works show that parameter-efficient fine-tuning performance (e.g., LoRA) can be worse than that of full fine-tuning, and that the gap in performance varies significantly across different tasks (Christophe et al., 2024a; Biderman et al., 2024). While our head-to-head comparisons based on parameter-efficient fine-tuning revealed that 7–9B-parameter medical models show limited improvement over their base models, the overall conclusions may be different in the full fine-tuning regime and at larger model scales.
While we acknowledge the limitations above, we do not believe they detract from the value of this work. We hope that our results call attention to a need for rigorous head-to-head evaluations when making similar claims of improved performance via medical DAPT, whether with other models, on other clinical tasks, or with respect to supervised fine-tuning versus zero-/few-shot prompting performance.
@@
We gratefully acknowledge DARPA (FA8750-23-2-1015), ONR (N00014-23-1-2368), NSF (IIS2211955), UPMC, Highmark Health, Abridge, Ford Research, Mozilla, the PwC Center, Amazon AI, JP Morgan Chase, the Block Center, the Center for Machine Learning and Health, and the CMU Software Engineering Institute (SEI) via Department of Defense contract FA8702-15-D-0002, for their generous support of our research.
@@ Appendix A Additional Details on Datasets
In this section, we provide additional details on the textual and visual medical QA datasets introduced in Section 3.
Table A1: Summary of the number of examples in the train, validation, and test sets for all textual medical knowledge QA (top), textual clinical note QA (middle), and visual QA datasets (bottom) used for evaluation.
@@
MMLU: Anatomy 5 14 135
MMLU: Clinical Knowledge 5 29 265
MMLU: College Biology 5 16 144
MMLU: College Medicine 5 22 173
MMLU: High School Biology 5 32 310
MMLU: Medical Genetics 5 11 100
MMLU: Nutrition 5 33 306
MMLU: Professional Medicine 5 31 272
MMLU: Virology 5 18 166
@@
i2b2 2008 Obesity Comorbidity Detection Challenge: Asthma 457 115 357
i2b2 2008 Obesity Comorbidity Detection Challenge: CAD 441 111 335
i2b2 2008 Obesity Comorbidity Detection Challenge: Diabetes 457 115 358
i2b2 2008 Obesity Comorbidity Detection Challenge: Obesity 443 111 334
CASI: Clinical Sense Disambiguation 9978 3304 3304
MIMIC-III: Clinical Sense Disambiguation 10384 3458 3458
Visual QA VQA-RAD 820 205 272
@@
MMMU: Basic Medical Science 5 5 25
MMMU: Clinical Medicine 5 5 25
MMMU: Diag. & Lab Medicine 5 5 25
MMMU: Pharmacy 5 5 25
MMMU: Public Health 5 5 25
In Table A1, we summarize the number of QA examples included in the train, validation, and test sets of each dataset, after following the preprocessing steps detailed in Section 3.
For VQA-RAD (Lau et al., 2018), PathVQA (He et al., 2020), and SLAKE (Liu et al., 2021), we only show the number of closed-ended visual QA examples, since our evaluations focus on closed-ended visual QA. For the datasets that required additional splits from the official train-validation-test split (e.g., due to the lack of a public test set), we include all of the fixed random seeds in our repository for reproducibility.
@@ Appendix B Additional Details on Model-Specific Prompt Selection
In this section, we provide additional details on how we define the prompt format search space discussed in Section 3.1. We construct a context-free grammar of plausible prompt formats following the approach by Sclar et al. (2024) (see Section 3.1 and Appendix A of the paper for reference). Using the Backus-Naur notation, we first define the basic fields $H_{q}$ for the question header (e.g., ‘‘### Question:’’), $H_{c}$ for the answer choice header (’e.g., ‘‘### Options:’’), and $H_{a}$ for the answer header (e.g., ‘‘### Answer:’’) as
@@
$L(f_{\text{wrap}},f_{\text{index}},n,s_{2})$ $::=C(f_{\text{wrap}},f_{\text{index}},0)s_{2}\ldots$
$\quad s_{2}C(f_{\text{wrap}},f_{\text{index}},n-1),$
@@
where we have left out the notations for the arguments for notational simplicity.
To define the prompt format search space, we instantiate the grammar above with the descriptors, separators, spaces, and functions shown below.
@@
Example 1: {mdframed} A key factor facilitating the application of nested case-control studies from the MACS was:
@@
Example 2: {mdframed} QUESTION – A key factor facilitating the application of nested case-control studies from the MACS was:
@@ Appendix C Additional Details on Zero-/Few-shot Prompting
Table C1: Summary of all of the prompting details made available for each medical LLM and VLM used for evaluation. For each column, a checkmark () indicates that the information was fully provided, a triangle () indicates that the information was partially provided (e.g., random sampling without information about the seeds), and a cross () indicates that the information is missing. “N/A” indicates that the corresponding information is not available due to its irrelevance to the evaluation setup considered in the paper (e.g., lack of few-shot example details because evaluation was only done in zero-shot or fine-tuning regimes).
Model System Prompt Zero-/Few-Shot Prompt Format Few-Shot fsexamples Sampling Details
Med42-v2 (Christophe et al., 2024b) N/A
Med42-v1 (Christophe et al., 2024a) N/A
OpenBioLLM (Pal and Sankarasubbu, 2024)
Clinical-Camel (Toma et al., 2023)
BioMistral (Labrak et al., 2024)
MediTron (Chen et al., 2023)
BioMedGPT-LM (Luo et al., 2023) N/A
LLaVA-Med (Li et al., 2023a) N/A
Med-Flamingo (Moor et al., 2023b)
@@ C.1 Reproducibility of Prompting Details
For the Med42-v2 models, we follow the instructions provided in the model cards posted by the authors on HuggingFace, for the 70B-parameter and 8B-parameter models. We use the recommended system prompt and the Llama-3-based conversational prompt format. Meanwhile, in Table C1, we treat the prompt format as partially missing, as the exact format that was used to format each question (“user” query) and answer (“assistant” response) for evaluation on closed-ended multiple-choice questions is not provided. As Christophe et al. (2024a) include the prompt format used for Med42-v1, we use this format since both models are from the same authors. In the original evaluation, the model predictions are selected based on the log-likelihood of each answer choice, which differs from our evaluation setup. We include the default prompt format used for Med42-v2 in Appendix C.2.1.
@@
For Med42-v1, we follow the instructions provided in the model card posted by the authors on HuggingFace. We use the recommended system prompt and the custom conversational prompt format with special <|system|>, <|prompter|>, and <|assistant|> tokens. For the prompt format, we refer to what is provided in Appendix A.1 of Christophe et al. (2024a). In the original evaluation, the model predictions are selected based on the log-likelihood of each answer choice, which differs from our evaluation setup. We include the default prompt format used for Med42-v2 in Appendix C.2.2.
@@
For the OpenBioLLM models, we follow the instructions provided in the model cards posted by the authors on HuggingFace, for the 70B-parameter and 8B-parameter models. We use the recommended system prompt and the Llama-3-based conversational prompt format. Meanwhile, in Table C1, we treat the prompt format as partially missing, as the exact format that was used to format each question (“user” query) and answer (“assistant” response) for evaluation on closed-ended multiple-choice questions is not provided. At the time of writing, there are no additional details about the models that have been publicly released, beyond what is provided in the model cards. We include the default prompt format used for OpenBioLLM in Appendix C.2.3.
@@
For Clinical-Camel, we use the conversational prompt format used in the official GitHub repository, which corresponds to the official chat format for Llama-2 (Touvron et al., 2023b). As the system prompts and few-shot examples used for the main evaluations in the paper are not provided, we use our own manually designed default system prompt and search over different choices of few-shot examples. For sampling, the evaluation code uses default temperature setting of 0.7 (albeit without the random seeds), which differs from our evaluation setup. We include the default prompt format used for Clinical-Camel in Appendix C.2.4.
@@
For BioMistral, we use the system prompt and zero-/few-shot prompt format provided in Appendix F of the paper. At the time of writing, the code repository is not publicly available, and the paper does not provide details on what few-shot examples were used for evaluation. In Section 4.3 of the paper, Labrak et al. (2024) mention that the output vocabulary is constrained to be one of the answer choices in lettered format (e.g., one of [A,B,C,D]) to force the model to avoid generating irrelevant tokens in its output. Meanwhile, it is not explicitly clear whether (i) the filtered token with the highest probability was treated as the model’s prediction or (ii) a token was randomly sampled based on the renormalized token probabilities. We also note that the vocabulary filtering procedure makes their evaluation setup different from ours, as we use greedy decoding to sample the model outputs without any constraints on the vocabulary (see Section 3). We include the default prompt format used for BioMistral in Appendix C.2.5.
@@
For the MediTron models, we use the system prompts—tailored specifically to MedQA, MedMCQA, PubMedQA, and the MMLU datasets—provided in Table 2 of the paper. For the prompt formats, we use the ones provided in the official GitHub repository, as the prompt formats (those with special ‘<|im_start|>’ and ‘<|im_end|>’ tokens, following the ChatML format) shown in the paper are only applicable to the fine-tuned models (see this discussion from the official GitHub repository). In particular, we refer to the prompt formats provided in the dataset preprocessing code and used for evaluation to determine the default prompt format for both the 70B- and 7B-parameter models. However, we were unable to reliably reproduce the zero-/few-shot prompting performance using this prompt format, and therefore perform a grid search over the prompt formats as well for model-specific prompt selection. In the evaluation code, Chen et al. (2023) provide the random seeds used for sampling the few-shot examples; however, we also search over the set of few-shot examples to consider a larger number of few-shot example choices. For sampling, we use the same greedy decoding approach as considered in the paper (referred to “Top Token Selection” in Section 4.3 of the paper). We include the default prompt format used for MediTron in Appendix C.2.6.
@@
While BioMedGPT-LM was evaluated on textual medical QA datasets such as MedMCQA and PubMedQA, the evaluation was performed only in the supervised fine-tuning regime, and the prompt formats used for these datasets are not available, to the best of our knowledge. Meanwhile, the official GitHub repository provides Jupyter notebook examples containing a conversational prompt format used in the context of other QA tasks. We therefore use this format by default but search over the prompt formats for model-specific prompt selection, since it is not specifically designed for closed-ended multiple-choice QA tasks. Moreover, as the system prompt provided is not semantically applicable to the QA tasks that we consider (e.g., ‘‘You are working as an excellent assistant in chemistry and molecule discovery.’’, we use our own manually designed default system prompt. We include the default prompt format used for BioMedGPT-LM in Appendix C.2.7.
@@
For LLaVA-Med, we use the system prompt and conversational prompt format included in the “simple_conv_med” template from the official GitHub repository (for LLaVA-v0 (Liu et al., 2023), we use the “simple_conv” template) by default. For formatting the visual questions, we refer to this file containing the raw visual QA results on VQA-RAD (‘‘Please choose from the following two options: [yes, no]’’). Meanwhile, we make these choices with the following caveats, to the best of our knowledge. First, the exact choice of system prompt and conversational prompt format used for evaluation are not discussed in the paper or the code repository, and we choose the one that has a system prompt specific to LLaVA-Med (‘‘You are LLaVA-Med, a large language and vision assistant trained by a group of researchers at Microsoft …’’) and follows the conversational format used for Vicuna-v0 (Chiang et al., 2023), which forms its LLM backbone. Second, details on how the answer choices should be formatted in the context of closed-ended QA tasks is only shown in the VQA-RAD results file. Given the uncertainty in such details, we also search over the prompt formats for model-specific prompt selection. We note that LLaVA-Med was not pretrained on multi-image inputs or evaluated in few-shot setting, and therefore details on the choice of few-shot examples are irrelevant. For sampling, the evaluation code uses a temperature setting of 0.7 (albeit without the random seeds), which differs from our evaluation setup. We include the default prompt format used for LLaVA-Med in Appendix C.3.1 and that for LLaVA-v0 in Appendix C.3.2.
@@
For Med-Flamingo, we use the system prompt and prompt format provided in the demo code from the official GitHub repository by default. However, we search over the prompt formats when performing model-specific prompt selection, as the example prompt in the demo does not show details for formatting answer choices in a closed-ended QA context. The choice of few-shot examples and the sampling details used for the original evaluations on VQA-RAD and PathVQA are not available. We include the default prompt format used for Med-Flamingo in Appendix C.3.3 and that for Open-Flamingo in Appendix C.3.4.

### v2

@@ 6 Discussion and Conclusion
In this work, we investigated the effectiveness of DAPT for training medically specialized LLMs and autoregressive VLMs suitable for various medical (visual) QA tasks. To that end, we compared several pairs of state-of-the-art medical LLMs/VLMs to their general-domain counterparts, whose only differences lie in medical DAPT and are exactly identical in model architecture and scale. Our work differs from prior works by providing direct apples-to-apples comparisons of medical and general-domain base models, while accounting for LLM/VLM sensitivity to prompting details and assessing the statistical significance of the results. We also ensure that our evaluations are comprehensive by (i) comparing the models in both the zero-/few-shot prompting and SFT regimes; and (ii) covering both QA tasks focused on assessing medical knowledge and those based on real-world clinical notes.
@@
In the zero-/few-shot prompting regime, we found that across both model classes and all model scales, the performance benefits from medical DAPT largely disappear when we (i) tailor the prompt format and choice of few-shot examples to each medical and general-domain model separately; and (ii) account for statistical uncertainty in model comparison (Section 4.1). In particular, we found that when we optimize the prompt only for the medical model and compare each model pair based on their absolute accuracies without accounting for uncertainty, the performance improvements from medical DAPT can be significantly overestimated (Section 4.2), potentially leading to unreliable conclusions about the benefits of medical DAPT on zero-/few-shot prompting performance.
@@
In the SFT regime, we found that medical LLMs overall do show statistically significant improvements on textual QA tasks focused on assessing medical knowledge (Section 5.1) but not on clinically relevant QA tasks that based on real-world clinical notes (Section 5.2). For medical VLMs, we found that they overall fail to show significant improvements on the visual medical QA tasks that we consider, which we present with the caveat that our findings are based on a relatively small number of VLMs and datasets (Section 5.3).
@@
Our findings suggest that for state-of-the-art general-domain LLMs and VLMs, the performance benefits from additionally pretraining on medical data from public sources such as PubMed may be limited. In fact, almost all of the medical models used in our evaluation use PubMed as the primary source of pretraining data for medical adaptation (Table 2), while open-source datasets commonly used for pretraining the general-domain base models in the first place (e.g., the Pile (Gao et al., 2020), S2ORC (Lo et al., 2020)) often already include PubMed data. Thus, we emphasize that any claims of improvement due to proposed medical DAPT procedures should be backed up by rigorous head-to-head comparisons against the corresponding general-domain base models.
Limitations.
@@
First, there is a vast and growing set of papers on applying medical DAPT to various general-domain base models, and we could not hope to compare all publicly available models here. While we selected the models to cover a wide range of general-domain base models and model scales (7B–70B) (Table 2) and included some of the latest models (e.g., OpenBioLLM and Llama-3), it is always possible that some newly released models do in fact yield better zero-/few-shot or SFT performance on medical QA.
@@
Third, due to computational constraints, we do not consider full fine-tuning of models for evaluations in the SFT regime. Prior works show that parameter-efficient fine-tuning performance (e.g., LoRA) can be worse than that of full fine-tuning, and that the gap in performance varies significantly across different tasks (Christophe et al., 2024a; Biderman et al., 2024). While our head-to-head comparisons based on parameter-efficient fine-tuning revealed that medical models show limited improvement over their base models across all model scales, the overall conclusions may be different in the full fine-tuning regime.
While we acknowledge the limitations above, we do not believe they detract from the value of this work. We hope that our results call attention to a need for rigorous head-to-head evaluations when making similar claims of improved performance via medical DAPT, whether with other models, on other clinical tasks, or with respect to SFT versus zero-/few-shot prompting performance.
@@
We gratefully acknowledge DARPA (FA8750-23-2-1015), ONR (N00014-23-1-2368), NSF (IIS2211955), UPMC, Highmark Health, Abridge, Ford Research, Mozilla, the PwC Center, Amazon AI, JP Morgan Chase, the Block Center, the Center for Machine Learning and Health, and the CMU Software Engineering Institute (SEI) via Department of Defense contract FA8702-15-D-0002, for their generous support of our research. We also thank Monica Agrawal, Ahmed Alaa, and Shan Chen for helpful discussions on an earlier version of this manuscript.
@@ Appendix A Additional Details on Datasets
Table A1: The numbers of train, validation, and test QA examples for all textual medical knowledge QA, textual clinical note QA, and visual medical QA datasets used for evaluation.
@@
MMLU (Anatomy) 5 14 135
MMLU (Clinical Knowledge) 5 29 265
MMLU (College Biology) 5 16 144
MMLU (College Medicine) 5 22 173
MMLU (High School Biology) 5 32 310
MMLU (Medical Genetics) 5 11 100
MMLU (Nutrition) 5 33 306
MMLU (Professional Medicine) 5 31 272
MMLU (Virology) 5 18 166
@@
i2b2 2008 Challenge (Asthma) 457 115 357
i2b2 2008 Challenge (CAD) 441 111 335
i2b2 2008 Challenge (Diabetes) 457 115 358
i2b2 2008 Challenge (Obesity) 443 111 334
CASI Sense Disambiguation 9978 3304 3304
MIMIC-III Sense Disambiguation 10384 3458 3458
Visual Medical QA VQA-RAD 820 205 272
@@
MMMU (Basic Medical Science) 5 5 25
MMMU (Clinical Medicine) 5 5 25
MMMU (Diag. & Lab Medicine) 5 5 25
MMMU (Pharmacy) 5 5 25
MMMU (Public Health) 5 5 25
For VQA-RAD (Lau et al., 2018), PathVQA (He et al., 2020), and SLAKE (Liu et al., 2021), we only show the number of closed-ended examples, since we focus on closed-ended QA. For the datasets that required additional splits from the official train-validation-test split (e.g., due to the lack of a public test set), we include all of the fixed random seeds.
@@ Appendix B Additional Details on Model-Specific Prompt Selection
To define the prompt format search space discussed in Section 3.1, we construct a context-free grammar of plausible prompt formats following Sclar et al. (2024) (see Section 3.1 and Appendix A of the paper for reference). Using the Backus-Naur notation, we first define the basic fields $H_{q}$ for the question header (e.g., ‘‘### Question:’’), $H_{c}$ for the answer choice header (’e.g., ‘‘### Options:’’), and $H_{a}$ for the answer header (e.g., ‘‘### Answer:’’) as
@@
$L(f_{\text{wrap}},f_{\text{index}},n,s_{2})$ $::=C(f_{\text{wrap}},f_{\text{index}},0)s_{2}\ldots s_{2}C(f_{\text{wrap}},f_{\text{index}},n-1),$
@@
where we exclude the arguments for simplicity. To define the prompt format search space, we instantiate the grammar with the descriptors, separators, spaces, and functions shown below.
@@
{custombox}
[frametitle=Example 1] A key factor facilitating the application of nested case-control studies from the MACS was:
@@
{custombox}
[frametitle=Example 2] QUESTION – A key factor facilitating the application of nested case-control studies from the MACS was:
@@ C.1 Reproducibility of Prompting Details
Table C1: Summary of prompting details shared for each medical LLM and VLM in the original papers. For each column, a ✓ indicates that the information was fully provided, and a ▲ indicates that the information was partially provided (e.g., random sampling without information about the seeds). Otherwise, the information was either not provided or irrelevant (e.g., no few-shot example details due to lack of few-shot evaluations).
Model System Prompt Prompt Format Few-Shot Examples Sampling Details
Med42-v2
Med42-v1
OpenBioLLM
Clinical-Camel
BioMistral
MediTron
BioMedGPT-LM
LLaVA-Med
Med-Flamingo
@@
We follow the instructions provided in the model cards on HuggingFace, for the 70B-parameter444huggingface.co/m42-health/Llama3-Med42-70B and 8B-parameter555huggingface.co/m42-health/Llama3-Med42-8B models. We use the recommended system prompt and the Llama-3-based conversational prompt format. Meanwhile, in Table C1, we treat the prompt format as partially missing, as the exact format that was used to format each question (“user” query) and answer (“assistant” response) for evaluation on closed-ended multiple-choice questions is not provided. As Christophe et al. (2024a) include the prompt format used for Med42-v1, we use this format since both models are from the same authors. In the original evaluation, the model predictions are selected based on the log-likelihood of each answer choice, which differs from our evaluation setup. We include the default prompt format used for Med42-v2 in Appendix C.2.1.
@@
We follow the instructions provided in the model card666huggingface.co/m42-health/med42-70b on HuggingFace. We use the recommended system prompt and the custom conversational prompt format with special <|system|>, <|prompter|>, and <|assistant|> tokens. For the prompt format, we refer to what is provided in Appendix A.1 of Christophe et al. (2024a). In the original evaluation, the model predictions are selected based on the log-likelihood of each answer choice, which differs from our evaluation setup (Section 3). We include the default prompt format used for Med42-v2 in Appendix C.2.2.
@@
We follow the instructions provided in the model cards on HuggingFace, for the 70B-parameter777huggingface.co/aaditya/Llama3-OpenBioLLM-70B and 8B-parameter888huggingface.co/aaditya/Llama3-OpenBioLLM-8B models. We use the recommended system prompt and the Llama-3-based conversational prompt format. In Table C1, we treat the prompt format as partially missing, as the exact format used to format each question (“user” query) and answer (“assistant” response) for closed-ended multiple-choice QA is not provided. At the time of writing, there are no additional details about the models that have been publicly released, beyond what is provided in the model cards. We include the default prompt format used for OpenBioLLM in Appendix C.2.3.
@@
We use the conversation format used in the official GitHub repository, which corresponds to that of Llama-2 (Touvron et al., 2023b). As the system prompts and few-shot examples used for the main evaluations in the paper are not provided, we use our own manually designed default system prompt and search over different choices of few-shot examples. For sampling, the evaluation code999github.com/bowang-lab/clinical-camel/blob/14d960f/evaluation/get_model_answer.py#L54 uses default temperature setting of 0.7 (albeit without the random seeds), which differs from our evaluation setup. We include the default prompt format used for Clinical-Camel in Appendix C.2.4.
@@
We use the system prompt and zero-/few-shot prompt format provided in Appendix F of the paper. At the time of writing, the code repository is not publicly available, and the few-shot example details are not fully disclosed. In Section 4.3 of Labrak et al. (2024), the authors mention that the output vocabulary is constrained to be one of the answer choices in lettered format (e.g., one of [“A”,“B”,“C”,“D”]) to force the model to avoid generating irrelevant tokens in its output. However, it is unclear whether (i) the filtered token with the highest probability was treated as the model’s prediction or (ii) a token was randomly sampled based on the re-normalized token probabilities. We include the default prompt format used for BioMistral in Appendix C.2.5.
@@
We use the system prompts—tailored specifically to MedQA, MedMCQA, PubMedQA, and the MMLU datasets—provided in Table 2 of the paper. For the prompt formats, we use the ones provided in the official GitHub repository, as the prompt formats (those with special ‘<|im_start|>’ and ‘<|im_end|>’ tokens, following the ChatML format101010 github.com/openai/openai-python/blob/release-v0.28.0/chatml.md) shown in the paper are only applicable to the fine-tuned models (see this discussion111111github.com/epfLLM/meditron/issues/13#issuecomment-1845955741 from the official GitHub repository). In particular, we refer to the prompt formats provided in the code121212github.com/epfLLM/meditron/blob/a7c7/evaluation/benchmarks.py#L622 code and used for evaluation131313github.com/epfLLM/meditron/blob/a7c7/evaluation/inference.py#L188 to determine the default prompt format for both the 70B- and 7B-parameter models. However, we were unable to reliably reproduce the zero-/few-shot prompting performance using this prompt format, and therefore perform a grid search over the prompt formats as well for model-specific prompt selection. In the evaluation code141414github.com/epfLLM/meditron/blob/a7c7/evaluation/inference.py#L188, Chen et al. (2023) provide the random seeds used for sampling the few-shot examples; however, we also search over the set of few-shot examples to consider a larger number of few-shot example choices. For sampling, we use the same greedy decoding approach as considered in the paper (“Top Token Selection” in Section 4.3 of the paper). We include the default prompt format used for MediTron in Appendix C.2.6.
@@
While BioMedGPT-LM was evaluated on medical knowledge QA tasks such as MedMCQA and PubMedQA, it was only evaluated in the SFT regime, and the prompt formats used for these datasets are not available to the best of our knowledge. Meanwhile, the official GitHub repository provides Jupyter notebook examples151515github.com/PharMolix/OpenBioMed/blob/main/examples/biomedgpt_inference.ipynb with a conversation format used in the context of other QA tasks. We thus use this format by default but search over the prompt formats for model-specific prompt selection, since it is not specifically designed for closed-ended multiple-choice QA tasks. Moreover, as the system prompt provided is not semantically applicable to the QA tasks that we consider (e.g., ‘‘You are working as an excellent assistant in chemistry and molecule di- scovery.’’, we use our own manually designed default system prompt. We include the default prompt format used for BioMedGPT-LM in Appendix C.2.7.
@@
For LLaVA-Med, we use the system prompt and conversational prompt format in the “simple_conv_med” template161616github.com/microsoft/LLaVA-Med/blob/b9a9/llava/conversation.py#L243 from the official GitHub repository (for LLaVA-v0 (Liu et al., 2023), we use the “simple_conv” template171717github.com/microsoft/LLaVA-Med/blob/b9a9/llava/conversation.py#L257) by default. For formatting the visual questions, we refer to this file181818github.com/microsoft/LLaVA-Med/blob/b9a9/llava/eval/eval_metrics/answer-file-llava-zeorshot.jsonl containing the raw visual QA results on VQA-RAD (‘‘Please choose from the following two options: [yes, no]’’). Meanwhile, we make these choices with the following caveats, to the best of our knowledge. First, the exact choice of system prompt and conversational prompt format used for evaluation are not discussed in the paper or the code repository, and we choose the one that has a system prompt specific to LLaVA-Med (‘‘You are LLaVA-Med, a large language and vision assistant trained by a group of researchers at Microsoft …’’) and follows the conversational format used for Vicuna-v0 (Chiang et al., 2023), which forms its LLM backbone. Second, details on how the answer choices should be formatted in the context of closed-ended QA tasks is only shown in the VQA-RAD results file. Given the uncertainty in such details, we also search over the prompt formats for model-specific prompt selection. We note that LLaVA-Med was not pretrained on multi-image inputs or evaluated in the few-shot setting, and therefore details on the choice of few-shot examples are irrelevant. For sampling, the evaluation code191919github.com/microsoft/LLaVA-Med/blob/b9a9/llava/eval/model_vqa_med.py uses a temperature setting of 0.7 (albeit without the random seeds), which differs from our evaluation setup. We include the default prompt formats used for LLaVA-Med and LLaVA-v0 in Appendices C.3.1C.3.2.
@@
We by default use the system prompt and prompt format provided in the demo code202020github.com/snap-stanford/med-flamingo/blob/7bcb/scripts/demo.py from the official GitHub repository. However, we search over the prompt formats when performing model-specific prompt selection, as the example prompt in the demo does not show details for formatting answer choices in a closed-ended QA context. The choice of few-shot examples and sampling details used for the original evaluations on VQA-RAD and PathVQA are not available. We include the default prompt formats used for Med-Flamingo and Open-Flamingo in Appendices C.3.3C.3.4.